Each of the equal sides of an isosceles triangle is 4 times the third side. The perimeter of the triangle is 144 inches. Find the sides of the triangle
Found 3 solutions by stanbon, richwmiller, dabanfield:
Answer by stanbon(75887)

Each of the equal sides of an isosceles triangle is 4 times the third side. The perimeter of the triangle is 144 inches. Find the sides of the triangle
Each of the equal sides of an isosceles triangle is 4 times the third side. The perimeter of the triangle is 144 inches. Find the sides of the triangle
----------
Let the 3third side be "x".
Then each equal side is "4x".
--------------
Equation:
x + 4x + 4x = 144
9x = 144
x = 16 inches (3rd side)
4x = 64 inches each of the equal sides.
======================================
Cheers,
Stan H.

Answer by richwmiller(17219)

x+4x+4x=144
9x=144
3*3x=3*4*3*4
x=16

Answer by dabanfield(803)

Each of the equal sides of an isosceles triangle is 4 times the third side. The perimeter of the triangle is 144 inches. Find the sides of the triangle
Let x be the length of the third side.
Then the other two "equal" sides are 4*x.
The perimeter then is:
x + 4*x + 4*x = 144
9*x = 144
x = 16
Calculate the other two sides substituting 16 for x.